Intercept coefficient- (1.05)

BadReports coefficient- (-1.36)

Income coefficient- (.026)

Choose the probability of CardAcceptance from the options below at the following values of predictors

BadReports=2 Income=50k

A)69.07%

B)30.93%

C)37%

D) 40.85%

it is logistic
Answer #1

Using the logistic model that is estimated, we can say that

the probability of Card Acceptance is

P(Card Acceptance = 1) = 21.05-1.36 Bad Report +0.026Income 1 + €1.05-1.36BadReport +0.026Income

where BadReports is the number of bad reports and Income is expressed in $1000s

Substituting BadReports=2 Income=50, we get

21.05-1.36 X 2+0.026x50 P(Card Acceptance = 1) = 1 1 + 1.05-1.36x2+0.026x50 = 0.4085

ans: the probability of CardAcceptance is D) 40.85%